1. A. The auditor sets AR to LOW, and RMM, the operating effectiveness, is found to be MODERATE. Calculate planned DR in accordance with the planning guideline. B. The auditor sets AR to LOW, and RMM, the operating effectiveness, is found to be HIGH. Calculate planned DR in accordance with the planning guideline. C. The auditor sets AR to MODERATE, and RMM, the operating effectiveness, is found to be LOW. Calculate planned DR in accordance with the planning guideline.
Answer #1

The Audit Risk at planning stage is governed by SA - 315

​Audit Risk means the risk that the auditor gives an inappropriate audit opinion when the financial statements are materially misstated. Thus, it is the risk that the auditor may fail to express an appropriate opinion in and audit assignment. Audit Risk is the function of RMM and DR :-    AR = RMM * DR Case A: AR = Low , RMM = Moderate Therefore, DR = AR/RMM,    Assume Low = 1, Moderate = 5, and High = 10

Therefore in case A DR = 1/5 = 0.20

Case B: AR = Low, RMM = High Therefore DR = 1/10 = 0.10    Case C: AR = Moderate, RMM = Low Therefore = 5/1 = 5 By this equation we can say that Higher the RMM lower will be the DR and Lower the RMM higher will be the DRR.

RMM may be defined as the risk that the financial statements are materially misstated prior to the audit. it consists of two elements :-

1. Inherent Risk = The suspectibility of an assertion about a class of transaction, account balance or disclosure to a misstatement that could be material, either individually or when aggregated with other misstatments, before consideration of any related controls.

2. Control Risk = The risk that a misstatement that could occur in an assertion about a class of transaction, account balance or disclosure and that could be material, either individually or when aggregated with other misstatements, will not be prevented, or detected and corrected, on a timely basis by the entity's internal control.

Detection Risk : The risk that the procedures performed by the auditor to reduce audit risk to an acceptably low level will not detect a misstatement that exists and that could be material, either individually or when aggregated with other misstatements.

RMM = IR*CR

AR = IR*CR*DR